Solve the following equation. Then place the correct answer in the box provided. Answer in terms of a mixed number.7(2x + 6) = 1
Answer by checkley79(3341)

7(2x + 6) = 1
14x+42=1
14x=1-42
14x=-41
x=-41/14 ans.
OR:
x=-2.9285714 ans.